Find the probability distribution for the given random variable. (Round all probabilities to four decimal places.) From a binwhat are the correct values for 0,1,2,3?
Answer #1

sol 27 27 50 = 0.4655 Po defective Joystics) 27 +31 p=0.4655 9 = 1- p = 1-0.4655 = 0.5345 3 И noa n q” P(x=0) Cx* p** 0.4655